# Sabiá-3 Technical Report

### v2

@@ 2.1 Multiple-choice Exams
We have developed a benchmark suite to evaluate the Sabiá-3 performance on 96 academic exams, including entry-level, undergraduate, and professional certification exams. Our focus is on multiple-choice questions from various Brazilian educational assessments, particularly those administered after the middle of 2023 to mitigate the risk of data contamination. However, ENADE is an exception, including spans of 2022 and 2023 to cover a wide range of disciplines.
For more details about the datasets, we recommend the readers to the Sabiá-2 technical report [5]. In addition to the 8 benchmarks introduced in [5], we now incorporated two new multiple-choice benchmarks:
@@
Table 2 shows the scores of the models in ten benchmarks derived from academic and professional exams applied after mid-2023. The results illustrate that Sabiá-3 has a comparable level of accuracy to GPT-4o across all 70 exams assessed. GPT-4o exhibits an advantage particularly in the Medicine benchmarks, while Sabiá-3 demonstrates a notable performance in the CPNU exams. Furthermore, Sabiá-3 surpasses Llama-3.1 405B by 3 percentage points in accuracy, performing equally or better in 63% of the exams. Claude 3.5 Sonnet achieves the highest average accuracy, although it presents a higher risk of data contamination due to its training data cutoff date.
@@
ENEM Multiple 2 83.8% 84.9% 83.8% 87.2% 89.9% 71.8% 81.0% 87.7%
BLUEX Multiple 2 77.3% 83.9% 85.1% 88.7% 93.0% 75.9% 79.5% 86.2%
ENADE Multiple 28 66.2% 73.0% 63.2% 75.2% 78.0% 64.6% 69.2% 77.2%
CPNU Multiple 8 82.2% 88.9% 83.4% 88.5% 83.8% 79.2% 84.4% 90.6%
BNDES Multiple 14 76.5% 80.0% 76.5% 86.7% 81.2% 75.5% 77.8% 84.3%
BCB Multiple 3 75.8% 79.4% 76.3% 84.1% 81.8% 65.0% 73.9% 80.1%
POSCOMP Comp. Sci 1 61.5% 78.5% 63.1% 73.8% 76.9% 50.8% 56.9% 76.9%
OAB Law 4 69.3% 77.4% 63.3% 83.1% 80.6% 72.1% 69.9% 75.9%
ENAM Law 2 50.6% 56.2% 50.6% 66.2% 75.0% 53.1% 53.1% 69.4%
CFCES Account. 1 50.0% 60.0% 60.0% 74.0% 76.0% 64.0% 54.0% 76.0%
Revalida Medicine 3 81.3% 84.2% 74.8% 87.1% 85.7% 73.0% 70.9% 83.5%
MREX Medicine 2 61.0% 61.5% 56.6% 68.7% 73.4% 49.9% 47.9% 57.2%
Average - 70 71.3% 77.0% 69.9% 80.6% 80.5% 68.9% 72.1% 80.2%
Table 2: Accuracy on academic benchmarks. The average results (last row) cover all the 70 exams applied after the middle of 2023, which is the knowledge cutoff date of Sabiá-3. Claude 3.5’s cutoff date is April of 2024, so exams released before that date might be present in its training data.
@@ 2.5 Function Calling Capabilities
GPT-4o 59.29% 85.52% 82.96% 71.79% 21.25% 63.41% 82.91%
GPT-4o Mini 58.45% 82.83% 81.80% 67.53% 25.75% 82.93% 71.83%
Sabiá-3 56.24% 88.75% 87.73% 68.33% 17.50% 97.56% 48.34%
Claude 3.5 Sonnet 54.83% 70.35% 66.34% 71.39% 23.50% 63.41% 75.91%
Sabiazinho-3 52.78% 88.88% 87.91% 65.57% 9.25% 78.05% 50.75%

### v3

@@ 2.1 Multiple-choice Exams
We have developed a benchmark suite to evaluate the Sabiá-3 performance on 93 academic exams, including entry-level, undergraduate, and professional certification exams. Our focus is on multiple-choice questions from various Brazilian educational assessments, particularly those administered after the middle of 2023 to mitigate the risk of data contamination. However, ENADE is an exception, including spans of 2022 and 2023 to cover a wide range of disciplines.
For more details about the datasets, we recommend the readers to the Sabiá-2 technical report [5]. In addition to the 8 benchmarks introduced in [5], we now incorporated five new multiple-choice benchmarks:
@@
  • BCB: The BCB benchmark refers to the public examination for the Central Bank of Brazil. This exam features objective questions with two possible answers: true or false. It comprises three distinct exams: basic knowledge, information technology, and economics and finance.
  • BNDES: The BNDES benchmark is the public examination for the National Bank for Economic and Social Development. Similar to the BCB exam, it consists of objective questions with two alternatives: true or false. The benchmark includes 14 exams covering a range of subjects such as law, systems analysis, administration, architecture, economics, engineering, accounting, and psychology.
  • CFCEQ: The CFCEQ benchmark is the technical qualification exam conducted by the Federal Accounting Council (CFC). Its purpose is to assess the knowledge and professional competence of accountants who wish to work as independent auditors in organizations regulated by the CVM, BCB, Susep, and Previc. The benchmark comprises 12 exams.
@@
Table 2 shows the scores of the models in 13 benchmarks derived from academic and professional exams applied after mid-2023. The results illustrate that Sabiá-3 has a comparable level of accuracy to GPT-4o across all 93 exams assessed. GPT-4o exhibits an advantage particularly in the Medicine benchmarks, while Sabiá-3 demonstrates a notable performance in the CPNU exams. Furthermore, Sabiá-3 surpasses Llama-3.1 405B by 2 percentage points in accuracy, performing equally or better in 62% of the exams. Claude 3.5 Sonnet achieves the highest average accuracy, although it presents a higher risk of data contamination due to its training data cutoff date.
@@
ENEM Multiple 2 83.8% 84.9% 83.8% 87.2% 89.9% 71.8% 82.7% 87.7%
BLUEX Multiple 2 77.3% 83.9% 85.1% 88.7% 93.0% 75.9% 83.7% 86.2%
ENADE Multiple 28 66.2% 73.0% 63.2% 75.2% 78.0% 64.6% 71.9% 77.2%
CPNU Multiple 8 82.2% 88.9% 83.4% 88.5% 83.8% 79.2% 85.5% 90.6%
BNDES Multiple 14 76.5% 80.0% 76.5% 86.7% 81.2% 75.5% 80.6% 84.3%
BCB Multiple 3 75.8% 79.4% 76.3% 84.1% 81.8% 65.0% 76.5% 80.1%
POSCOMP Comp. Sci 1 61.5% 78.5% 63.1% 73.8% 76.9% 50.8% 76.9% 76.9%
OAB Law 5 68.6% 77.6% 64.1% 83.4% 81.9% 72.1% 72.6% 76.4%
ENAM Law 3 53.8% 59.3% 50.0% 70.2% 75.2% 58.9% 62.6% 70.6%
CFCES Account. 4 61.3% 68.9% 65.3% 75.4% 80.4% 59.3% 70.4% 78.4%
CFCEQ Account. 12 61.3% 72.1% 58.3% 73.5% 79.2% 63.0% 66.7% 70.7%
Revalida Medicine 3 81.3% 84.2% 74.8% 87.1% 85.7% 73.0% 76.2% 83.5%
MREX Medicine 8 65.5% 69.0% 60.9% 76.1% 73.7% 53.4% 52.4% 63.7%
Average - 93 70.4% 76.9% 69.6% 80.8% 81.6% 66.4% 73.7% 79.0%
Table 2: Accuracy on academic benchmarks. The average results (last row) cover all the 93 exams applied after the middle of 2023, which is the knowledge cutoff date of Sabiá-3. Claude 3.5’s cutoff date is April of 2024, so exams released before that date might be present in its training data.
@@ 2.5 Function Calling Capabilities
GPT-4o 69.58% 87.42% 89.02% 79.65% 41.00% 83.33% 83.15%
GPT-4o Mini 64.10% 85.21% 83.57% 74.41% 34.12% 83.33% 74.75%
Sabiá-3 56.50% 90.65% 91.46% 75.83% 7.00% 88.89% 59.90%
Claude 3.5 Sonnet 56.46% 45.44% 47.89% 78.94% 41.00% 77.78% 74.04%
Sabiazinho-3 55.21% 86.62% 82.67% 75.08% 7.25% 77.78% 74.61%